3/4𝑥+3≤2/3≤2/𝑥+5


1
Expert's answer
2021-05-16T19:11:02-0400

"{3\over 4x} +3\leq {2\over 3}\leq {2\over x}+5"

Split into two and solve separately

1). "{3\over 4x}+3\leq {2\over 3}"

Multiply through by 4x

"{3\over 4x}(4x)+3(4x)\leq {2(4x)\over 3}"

"\implies 3+12x\leq {8x\over 3}"

Multiply through by 3

"3(3)+12x(3)\leq{8x\over 3}(3)"

"\implies 9+36x\leq8x"

"9-9+36x-8x\leq8x-8x-9"

"\implies 28x\leq -9\implies x\leq {-9\over 28}"

2). "{2\over 3}\leq{2\over x}+5"

Multiply through by 3x

"{2\over 3}(3x)\leq{2\over x}(3x)+5(3x)"

"\implies 2x\leq 6+15x \implies 2x-2x-6\leq6-6+15x-24"

"\implies -6\leq 13 x \implies x\geq {-6\over 13}"

Combining the two,we have

"{-6\over 13}\leq x\leq {-9\over 28}"